OpenStudy (anonymous):

Find the exact circumference of a circle with an area equal to 36 sq. in.

OpenStudy (dls):

\[C=2\Pi R = 36 \] From here, R=5.73 Area=pi*r*r=>18*3.14=56.52 sq in.

OpenStudy (dls):

Did you get it @mjenkins ?

OpenStudy (lgbasallote):

i think you got your thingies criss-crossed @DLS area = 36 find the circumference

OpenStudy (dls):

oh lol ..sorry Area=36 pi*r*r=36 r*r=11.14 r=3.38 C=2piR =>2*3.14*3.38=21.26sq in.

OpenStudy (anonymous):

thank you !

OpenStudy (dls):

np!

OpenStudy (amistre64):

exact values that use pi generally do not stop at 2 decies

OpenStudy (amistre64):

but yes A = pi r^2 sqrt(A/pi) = r C = 2r pi C = 2sqrt(A/pi) * pi

OpenStudy (anonymous):

can one of you tell me how to ask another question ? lol

OpenStudy (dls):

Close this one

OpenStudy (theviper):

There is an option on your question \(\Large{Close}\). Just click & ask another :)

OpenStudy (anonymous):

ok . thanks